OpenStudy (anonymous):

Help me solve this question please: onefourthy + 9 = onehalf

OpenStudy (rulnick):

(1/4)y + 9 = 1/2 (1/4)y = -8.5 y = -34

OpenStudy (anonymous):

Thank you, you where a big help :)